what equation is is used; "put equation in slope-intercept form by solving for y" problem is 3y = 4x
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

what equation is is used; "put equation in slope-intercept form by solving for y" problem is 3y = 4x
----
y" normally designates the 2nd derivative.
If that is what you mean,
If y = (4/3)x
then y' = 4/3
and y" = 0
================
Cheers,
Stan H.
================

Answer by Alan3354(69443)

"put equation in slope-intercept form by solving for y" problem is 3y = 4x
-----------
3y = 4x
y = 4x/3
or y = (4/3)x + 0